OpenStudy (anonymous):

Calculus? Use implicit diferentiation to find dy/dx for sin x^3 = cos y ^2 Please help?

OpenStudy (btaylor):

The derivative of the left side (with respect to x -- d/dx) is 3x^2 cos(x^3). The derivative of the right side (again, with respect to x) is -2y sin(y^2)(dy/dx). Solving for dy/dx yields:\[=\frac{3x^2\cos(x^3)}{-2y \sin(y^2)}\]

OpenStudy (anonymous):

great thank you :)

OpenStudy (btaylor):

yw

OpenStudy (anonymous):

could you help me on one more?

OpenStudy (anonymous):

how do i solve this?

OpenStudy (anonymous):

@BTaylor please tell me, can you help?